# Is This Collection Worth My LLM’s Time? Automatically Measuring Information Potential in Text Corpora

### v2

@@ Abstract
As large language models (LLMs) converge towards similar capabilities, the key to advancing their performance lies in identifying and incorporating valuable new information sources. However, evaluating which text collections are worth the substantial investment required for digitization, preprocessing, and integration into LLM systems remains a significant challenge. We present a novel approach to this challenge: an automated pipeline that evaluates the potential information gain from text collections without requiring model training or fine-tuning. Our method generates multiple choice questions (MCQs) from texts and measures an LLM’s performance both with and without access to the source material. The performance gap between these conditions serves as a proxy for the collection’s information potential. We validate our approach using five strategically selected datasets: EPFL PhD manuscripts, a private collection of Venetian historical records, two sets of Wikipedia articles on related topics, and a synthetic baseline dataset. Our results demonstrate that this method effectively identifies collections containing valuable novel information, providing a practical tool for prioritizing data acquisition and integration efforts.
@@ 1 Introduction
Recent advances in large language models (LLMs) have revealed a striking phenomenon: as these models scale, they tend to develop remarkably similar internal representations and capabilities regardless of their architecture Huh et al. (2024). This convergence, coupled with established scaling laws Kaplan et al. (2020) and the growing recognition that AI development should shift from model-centric to data-centric approaches Zha et al. (2025), suggests that the key to advancing LLM capabilities lies in identifying and incorporating novel, high-quality information sources rather than architectural innovations. However, identifying valuable text collections for model enhancement presents significant challenges: digitization and preprocessing are costly, and training or fine-tuning models on new data requires substantial computational resources.
This creates a critical need: how can we efficiently evaluate whether a text collection contains information that would meaningfully expand an LLM’s knowledge? Current approaches typically require actually training or fine-tuning models on new data to assess its value - an expensive and time-consuming process. While retrieval-augmented generation (RAG) Lewis et al. (2020) offers a promising approach for helping models access long-tail knowledge without full retraining Kandpal et al. (2023), this still requires careful curation of knowledge bases and comes with its own computational costs. Moreover, the challenge of identifying valuable text collections remains: digitization and preprocessing are costly, and organizations need ways to evaluate the information potential of document collections before investing in their integration.
We present a novel approach to this challenge: an automated pipeline that evaluates the potential information gain from text collections without requiring model training or fine-tuning. Our method generates and leverages multiple choice questions (MCQs) to systematically probe whether the knowledge contained in a text collection is already accessible to an LLM or represents genuinely novel information.
To rigorously validate our approach, we evaluate it across three strategically selected datasets: (1) EPFL PhD manuscripts, containing specialized academic knowledge likely novel to LLMs, (2) Wikipedia articles that were presumably part of the LLMs’ training data, and (3) a synthetic baseline dataset composed of model-generated texts. Our key contributions are:
  1. 1.
    A novel end-to-end pipeline that uses MCQs to efficiently evaluate the potential information gain from text collections without requiring model training or fine-tuning
  2. 2.
    An automated filtering methodology for ensuring MCQ quality through complementary similarity metrics that address two key challenges: verifying question relevance and grounding in source material, while ensuring distractors are plausible yet unambiguously incorrect
  3. 3.
    A systematic analysis of knowledge gaps across diverse text collections that identifies areas of model uncertainty, quantifies the potential value of different information sources, and guides strategic decisions about data collection and integration
Importantly, our approach is dataset-agnostic and can be applied to any text corpus. This makes it applicable to different domains and research questions. This pipeline could be leveraged to select textual data for model enhancements through fine-tuning or retrieval-augmented generation systems.
@@ 2 Related Work
Knowledge Detection in LLMs. Prior research has developed several methods to analyze how LLMs process and retain textual information. Work on memorization Hartmann et al. (2023); Shi et al. (2024b) and data contamination Yax et al. (2024); Golchin and Surdeanu (2024) focuses on identifying verbatim recall of training data, while hallucination detection Farquhar et al. (2024) aims to identify when models generate false information. Research on novelty detection has primarily focused on linguistic and semantic novelty McCoy et al. (2023); Lin et al. (2024), with less attention paid to factual novelty. While these approaches provide valuable insights into model behavior, they are retrospective – analyzing what models have already learned or memorized. In contrast, our work takes a prospective approach, developing metrics to evaluate the potential value of new information sources before investing in their integration into LLM systems.
@@ 3 Methods
Multiple Choice Questions (MCQs) are a well-established tool for knowledge assessment, supported by research in cognitive science and educational psychology Thomas M. Haladyna and Rodriguez (2002). Their four-option format, consisting of one correct answer and three distractors, offers an optimal balance between assessment reliability and cognitive load Vyas and Supe (2007). MCQs are particularly valuable for automated evaluation as they provide objective correctness measures while efficiently testing understanding across diverse topics Oc and Hassen (2024). When working with Large Language Models (LLMs), MCQs offer an additional advantage: they constrain the output space to a finite set of options, eliminating the ambiguity and variability inherent in open-ended responses and enabling precise evaluation of model knowledge.
Generating high-quality MCQs presents unique challenges, particularly due to the absence of ground truth signals typically available in human-curated educational assessments. Without explicit supervision on what constitutes a good question or appropriate distractors, we must derive reliable metrics to filter and validate the generated MCQs. This automated filtering setting raises two critical challenges: ensuring questions are both relevant and non-trivial, and guaranteeing that distractors are plausible yet unambiguously incorrect. The latter is particularly crucial when evaluating LLMs – if distractors are too simple, models might succeed through elimination rather than true knowledge; if too similar to the correct answer, they might confuse models even when provided with context. To address these challenges, we propose similarity-based filtering metrics that serve as proxy signals for question quality and difficulty, enabling automated quality control without human supervision.
Figure 1: Overview of our knowledge evaluation pipeline. The framework consists of three main components: (1) MCQ Generation using LLMs to create questions from input datasets, (2) Quality-focused filtering using similarity metrics, and (3) Position-debiased evaluation to assess model knowledge with and without context. ($c$: context, $q$: question, $g$: ground truth, $d_{i}$: distractors)
3.1 Pipeline
As illustrated in Figure 1, our pipeline consists of three key stages: (1) MCQ generation from text using LLMs, (2) quality-focused filtering using similarity metrics, and (3) position-debiased evaluation to assess model knowledge both with and without context.
3.2 MCQ Generation and Filtering
Question Generation Process. The input text is first divided into manageable chunks of 2000 words to ensure consistent context length across questions. These chunks serve as the basis for LLM-generated MCQs, where each question is crafted to test understanding of specific information within the chunk. The exact prompts used for generation can be found in table 2 of the Supplementary Methods.
Two-Stage Quality Filtering. To address the dual challenges of ensuring question relevance and appropriate distractor difficulty without human supervision, we employ two complementary filtering mechanisms:
@@ 3.2 MCQ Generation and Filtering
    Context-Answer Alignment Filter: To ensure questions are both relevant and grounded in the source material, we verify that the correct answer is more strongly aligned with the source context than any distractor:
    $\min_{i}\left[\text{sim}(c,g)-\text{sim}(c,d_{i})\right]\quad\text{for}\ i=1,2,3$ (1)
    where similarity is measured using both Jaccard index and ROUGE-L score Lin (2004). This helps eliminate misleading or incorrect MCQs while ensuring questions test information actually present in the context.
@@
    Distractor Plausibility Filter: To prevent trivial questions while maintaining unambiguous correctness, we ensure distractors are semantically similar to the correct answer but not identical:
    $\max_{i}\left[\text{cos-sim}(g,d_{i})\right]\quad\text{for}\ i=1,2,3$ (2)
    using NVIDIA’s state-of-the-art NV-Embed-v2 model de Souza P. Moreira et al. (2024); Lee et al. (2024) for text embeddings. This creates challenging questions where distractors are plausible enough to require true understanding while remaining distinctly incorrect.
3.3 Evaluation
We evaluate the information potential ($IP$) of a context $c$ for a given LLM $f$ with Eq 3 as the improvement in correctness when incorporating context in the prompt. Formally, it is defined as:
$IP=\frac{C_{\text{context}}-C_{\text{direct}}}{|\mathcal{Q}|-(I_{\text{context}}+I_{\text{direct}})}$ (3)
where:
$C_{\text{context}}$ $=\sum_{q,c\in\mathcal{Q}}\mathbbm{1}(f(q|c)=g),\quad C_{\text{direct}}=\sum_{q\in\mathcal{Q}}\mathbbm{1}(f(q)=g),$
$I_{\text{context}}$ $=\sum_{q,c\in\mathcal{Q}}\mathbbm{1}(f(q|c)\neq g),\quad I_{\text{direct}}=\sum_{q\in\mathcal{Q}}\mathbbm{1}(f(q)\neq g).$
Here, $\mathbbm{1}(\cdot)$ is an indicator function equal to 1 if the condition is true, and 0 otherwise. The denominator excludes questions that the model gets wrong both with and without context, ensuring we only measure improvement on questions where success is possible in at least one condition.
Position Bias Mitigation. To address the known issue of positional bias in LLMs Shi et al. (2024a), we evaluate each MCQ four times, rotating the correct answer through all possible positions (A, B, C, D) while randomly arranging the distractors. This effect can be seen in Figures 4, 5, and 6. This ensures that model performance reflects true knowledge rather than position-based preferences.
@@ 3.4 Data
To evaluate our pipeline’s efficacy, we conduct experiments across five strategically selected datasets representing distinct knowledge domains: (1) 177 EPFL PhD manuscripts containing specialized academic research, (2) a private collections of historical records on Venetian urban history, (3) Wikipedia articles related to EPFL manuscript topics, (4) Wikipedia articles connected to Venetian history, and (5) synthetic LLM-generated text. This dataset selection facilitates a systematic evaluation of knowledge coverage and novelty.
We hypothesize that PhD manuscripts and the Venetian historical records will yield the highest knowledge gaps, as they both contain information likely absent from training data - recent research contributions in the case of PhD manuscripts, and rare historical records in the case of the Venetian collection. The synthetic dataset, comprising model-generated content, serves as a lower-bound baseline for knowledge novelty. Wikipedia articles, being a primary source for LLM training Brown et al. (2020); Touvron et al. (2023), function as a control group representing knowledge presumably well-embedded in the models’ parameters.
Data collection. The EPFL PhD manuscripts were systematically acquired from the university’s institutional repository, which functions as the authoritative digital archive for doctoral dissertations.
The Venetian urban history collection consists of public domain books about Venice and its history. The vast majority had not been digitized previously, suggesting that they contain historical knowledge likely absent from the training data of contemporary large language models (LLMs).
The Wikipedia datasets are created by leveraging the Wikipedia API. With the help of the PhD manuscript titles and Venetian history topics, related articles are fetched and combined into similar-sized texts. Similarly, these titles and topics are used to generate the baseline dataset. To overcome the LLMs’ generation limits, in a first step, subtopics are generated surrounding the manuscript titles and Venetian history themes. Then, chunks of around 600 words are generated, using the subtopics as overarching themes, and concatenated.
All data used in our experiments (including chunks, synthetic and baseline datasets) as well as the complete code of our evaluation pipeline are made available open-source with this submission under creative commons license.
@@ 4 Results
GPT-4o Wikipedia-EPFLLlama 70B BaselineLlama 70B Wikipedia-EPFLGPT-4o Wikipedia-VeniceGPT-4o EPFLLlama 70B EPFLGPT-4o Venice000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Fraction of Correct ResponsesIP ScoreBoth CorrectWith ContextNo ContextBoth Incorrect
@@
Our analysis reveals three key findings: (1) similarity thresholding effectively increases question difficulty while maintaining answerable questions, (2) PhD manuscripts and Venetian records contain significantly more novel information compared to Wikipedia and synthetic datasets, and (3) larger language models show consistently lower information potential across all remaining datasets, suggesting better knowledge retention during pre-training. We present detailed evidence for each of these findings below.
4.1 Effectiveness of Similarity Thresholding
001010101020202020303030304040404050505050606060600.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Cutoff Percentile ValuesFraction of Correct ResponsesCosineRouge-L and JaccardWC 4xNC 4xQuestions Remaining
Effect of Cosine Similarity Thresholding. Figure 3 displays the effect of cosine thresholding. It shows a strong decrease of around 10% between the absence of thresholding and the 50th percentile cutoff in the performance of the model with no context (NC 4x), while that of the model with context (WC 4x) remains steady. Therefore, the cosine similarity has the expected effect of increasing the MCQ difficulty when answering without context while maintaining the context-based performance.
Effect of Jaccard and ROUGE-L Thresholding. A different trend is observed when applying the Jaccard and ROUGE-L thresholding, as seen in Figure 3. Here, a slight increase in performance with context (WC 4x) of 2% can be observed between no thresholding and the 50th percentile cutoff, while performance without context (NC 4x) remains stable.
Comparative Analysis of Thresholding Methods. Our results demonstrate that cosine similarity thresholding is effective at increasing question difficulty while Jaccard and ROUGE-L slightly improve question clarity shown by the small increase in performance when the model is extended with context. Due to the separate percentile-based filtering of Jaccard and ROUGE-L scores followed by their combination, a substantial drop in the number of retained questions is observed between the 50th and 60th percentiles.
4.2 Information Potential (IP) Across Different Datasets
Limited-Access Knowledge Collections (IP: 0.229 - 0.265). Both the EPFL PhD manuscripts and Venetian historical records yield substantially higher information potential than other tested collections. We observe significant performance gaps between context-free (73.4% and 70.5%, respectively) and context-provided (98% and 97.5%) conditions, indicating considerable novel information content. These complementary results confirm our hypothesis that specialized collections with limited digital accessibility—whether recent academic research or rare historical records—represent particularly valuable information sources for LLMs, containing knowledge that is hard to retrieve or not well-represented in their training data.
Wikipedia Datasets (IP: 0.110 - 0.180). Both Wikipedia datasets demonstrate moderate information potential, with notably smaller gaps between context-free and context-provided performance (83.6% vs. 97.8% for EPFL-related Wikipedia articles and 82.8% vs. 97.2% for Venetian history Wikipedia articles). This attenuated differential suggests that a substantial portion of this information is already encoded in the models’ parameters. We hypothesize that these Wikipedia articles, or semantically equivalent content, were included in the models’ training corpora, allowing them to answer derived MCQs with relatively high accuracy even without contextual assistance. This observation corroborates our initial hypothesis and validates our methodological approach of using Wikipedia as a controlled baseline.
Synthetic Baseline Dataset (IP: 0.125). Among the three datasets, the synthetic baseline yields the lowest information potential. We observe a minimal performance gap between conditions (85.5% vs. 98.8%), with high performance in the context-free setting reflecting the model’s inherent familiarity with self-generated content. This dataset effectively serves as a lower bound for the information potential metric, providing a useful reference point for evaluating other collections.
@@ 4.3 Comparison between open and closed source model
Despite their architectural and size differences, both Llama 70B and GPT-4o exhibit similar patterns in information potential across datasets, though with notable variations in magnitude. GPT-4o shows slightly lower information potential scores compared to Llama 70B, suggesting better baseline knowledge retention in the larger model. Interestingly, the relative gap between datasets remains consistent across both models - EPFL manuscripts consistently show approximately double the information potential compared to Wikipedia. This consistency across different model architectures and training approaches strengthens the validity of our information potential metric. Additionally, both models maintain near-perfect performance (>97%) when provided with context, regardless of the dataset, indicating robust comprehension capabilities when given access to relevant information.
While a more comprehensive analysis comparing a wider range of LLMs could provide additional insights into the relationship between model architecture, size, and information potential, such an investigation falls outside the scope of this study, which focuses on establishing the validity of our evaluation methodology.
4.4 Qualitative Analysis of High-Value Information
MCQ (correct answer in italics) Relevant Context Passages
B) Both typically result from incorrect or incomplete mental models.
C) Both involve action reversals.
A) Repetition.
B) Omission.
C) Timing error.
D) Sequence error.
A) AFs rely on spectral analysis.
B) AFs use a frame-to-phoneme alignment.
C) AFs map phonemes to articulatory features.
D) AFs predict phonemes directly.

### v3

@@ Abstract
As large language models (LLMs) converge in capabilities, progress increasingly depends on integrating novel, high-value information sources. Yet deciding which text collections justify the costs of digitization and integration remains difficult. We introduce an automated pipeline that estimates a collection’s information potential without retraining. The method generates multiple-choice questions (MCQs) from texts and compares model performance with and without access to the source; the gap serves as a proxy for novel knowledge. We validate this approach on five datasets—EPFL PhD manuscripts, Venetian historical records, two Wikipedia subsets, and a synthetic baseline. Results show that the pipeline reliably distinguishes high-value from redundant corpora, offering a practical tool for prioritizing data acquisition and integration.
@@ 1 Introduction
Large language models (LLMs) show convergent capabilities as they scale, largely independent of architecture Huh et al. (2024). Combined with scaling laws Kaplan et al. (2020) and the shift toward data-centric AI Zha et al. (2025), this suggests that future progress depends less on model design and more on integrating high-quality, novel information sources. Yet acquiring such data is costly: digitization, preprocessing, and fine-tuning require substantial resources.
This raises a central question: how can we assess whether a text collection meaningfully expands an LLM’s knowledge before investing in its integration? Existing strategies typically require expensive retraining or fine-tuning, while retrieval-augmented generation (RAG) Lewis et al. (2020) mitigates some costs but still requires careful knowledge curation Kandpal et al. (2023). Practical methods for estimating the value of new corpora remain limited.
We propose an automated pipeline that estimates a collection’s information potential without training. The method generates MCQs from text chunks, applies a two-stage quality filter—(i) context–answer alignment using ROUGE-L/Jaccard and (ii) distractor plausibility via embedding cosine similarity—and then compares LLM performance with and without access to the source. The performance gap serves as a proxy for novel information.
To validate this approach, we test three distinct datasets: (1) EPFL PhD manuscripts, containing specialized academic knowledge, (2) Wikipedia articles, representing widely available content, and (3) a synthetic baseline of model-generated text. Importantly, the pipeline is dataset-agnostic and can be applied to any domain, making it a practical tool for prioritizing corpora for fine-tuning or RAG systems.
@@ 2 Related Work
Knowledge Detection in LLMs. Prior research has developed several methods to analyze how LLMs process and retain textual information. Work on memorization Hartmann et al. (2023); Shi et al. (2024) and data contamination Yax et al. (2024); Golchin and Surdeanu (2024) focuses on identifying verbatim recall of training data, while hallucination detection Farquhar et al. (2024) aims to identify when models generate false information. Research on novelty detection has primarily focused on linguistic and semantic novelty McCoy et al. (2023); Lin et al. (2024), with less attention paid to factual novelty. While these approaches provide valuable insights into model behavior, they are retrospective – analyzing what models have already learned or memorized. In contrast, our work takes a prospective approach, developing metrics to evaluate the potential value of new information sources before investing in their integration into LLM systems.
@@ 3 Methods
Multiple Choice Questions (MCQs) are a well-established tool for knowledge assessment, supported by research in cognitive science and educational psychology Thomas M. Haladyna and Rodriguez (2002). Their four-option format, consisting of one correct answer and three distractors, offers an optimal balance between assessment reliability and cognitive load Vyas and Supe (2007). MCQs are particularly valuable for automated evaluation as they provide objective correctness measures while efficiently testing understanding across diverse topics Oc and Hassen (2024). For LLM evaluation, they offer the added advantage of constraining the output space to a finite set of options, eliminating ambiguity in grading and enabling precise comparisons across conditions.
3.1 MCQ Generation
The input text is first divided into manageable chunks of 2000 words to ensure consistent context length across questions. These chunks serve as the basis for LLM-generated MCQs, where each question is crafted to test understanding of specific information within the chunk. The exact prompts used for generation can be found in table 1 of the Supplementary Methods.
3.2 MCQ Quality Filtering (Two-Stage)
To avoid trivial or ambiguous items, we apply complementary filters that test (a) grounding in the source and (b) distractor plausibility:
(a) Context–Answer Alignment.
For each MCQ, we require the correct answer $g$ to be better aligned with its source chunk $c$ than any distractor $d_{i}$. We compute surface-level similarity via Jaccard and ROUGE-L Lin (2004):
$\min_{i}\left[\text{sim}(c,g)-\text{sim}(c,d_{i})\right]\quad\text{,}\ i=1,2,3$ (1)
where $\text{sim}\in\{\text{Jaccard},\text{ROUGE-L}\}$. Items failing either metric are discarded, removing questions whose “correct” answer is not actually supported by the text, or whose distractors inadvertently match the context.
(b) Distractor Plausibility.
To prevent elimination-by-obviousness while keeping answers unambiguous, we enforce semantic closeness (but not identity) between $g$ and at least one distractor using cosine similarity over sentence embeddings using the state-of-the-art NV-Embed-v2 de Souza P. Moreira et al. (2024); Lee et al. (2024):
$\max_{i}\left[\text{cos-sim}(g,d_{i})\right]\quad\text{,}\ i=1,2,3$ (2)
which retains items where distractors are “nearby” enough to be plausible but not paraphrases of the correct answer. Thresholds are set by percentiles (e.g., 40–60th) computed on the pool of generated items, trading coverage for quality.
3.3 Evaluation (Position-Debiased, With/Without Context)
Each filtered MCQ is evaluated in two conditions:
@@ 3.3 Evaluation (Position-Debiased, With/Without Context)
    Direct (no context): The model answers $(q,\{A,B,C,D\})$ without access to the chunk.
@@
    With context: The model receives the chunk $c$ plus the same $(q,\{A,B,C,D\})$.
To mitigate position bias, we evaluate each item four times, rotating the correct option through A, B, C, and D while shuffling distractors. A question is marked correct in a condition only if the model selects the correct option for all four rotations (a conservative criterion that reduces spurious gains).
We define the information potential ($IP$) of a collection as the normalized improvement in correctness when adding context:
$IP=\frac{C_{\text{context}}-C_{\text{direct}}}{|\mathcal{Q}|-(I_{\text{context}}+I_{\text{direct}})},$ (3)
where $C_{\text{context}}$ and $C_{\text{direct}}$ count questions answered correctly (with the “4x” position control) with and without context, and $I_{\text{context}},I_{\text{direct}}$ count incorrect ones. The denominator excludes items missed in both settings to focus on questions that are answerable in at least one condition.
@@ 3.4 Data
We evaluate five collections with distinct expected novelty profiles:
  • EPFL PhD manuscripts: 177 theses with specialized, recent research likely underrepresented in pretraining corpora.
  • Venetian historical records: Public-domain books on Venice’s urban history, many newly digitized and rare online.
  • Wikipedia (EPFL-related) and Wikipedia (Venice-related): Articles fetched via the Wikipedia API using manuscript titles and Venetian themes as seeds; expected to overlap with pretraining data.
  • Synthetic baseline: Model-generated chunks on the same themes to establish a lower-bound novelty reference.
All corpora are chunked uniformly; MCQs, filtering metadata, and prompts are released with this submission to support reproducibility.
@@ 4 Results
Across five corpora and two models, our pipeline (i) increases MCQ quality via complementary similarity filters, (ii) assigns higher information potential (IP) to specialized/rare collections than to widely available ones, and (iii) yields consistent cross-model patterns. Figure 1 illustrates the impact of different thresholding methods, while Figure 2 summarizes correctness with/without context and the resulting IP per dataset/model.
4.1 Effectiveness of Similarity Thresholding
01010202030304040505060600.30.30.40.40.50.50.60.60.70.70.80.80.90.911Cutoff Percentile ValuesFraction of Correct ResponsesCosineRouge-L and JaccardWC 4xNC 4xQuestions Remaining
Cosine similarity. Increasing the cutoff reduces performance without context (NC 4x) by about 10% between 0 and the 50th percentile, while performance with context (WC 4x) stays stable—indicating higher difficulty for no-context answers without hurting context-based performance.
ROUGE-L and Jaccard. In contrast, these filters slightly improve WC 4x accuracy (+2% up to the 50th percentile) while NC 4x remains stable, suggesting improved question clarity rather than difficulty. However, combined percentile filtering sharply reduces the number of retained questions between the 50th and 60th percentiles.
Comparison. Cosine thresholding mainly increases difficulty, whereas ROUGE-L and Jaccard refine clarity at the cost of coverage.
4.2 Information Potential Across Corpora
GPT-4o Wikipedia-EPFLLlama 3 70B BaselineLlama 3 70B Wikipedia-EPFLGPT-4o Wikipedia-VeniceGPT-4o EPFLLlama 3 70B EPFLGPT-4o Venice00.20.20.40.40.60.60.80.811Fraction of Correct ResponsesIP ScoreBoth CorrectWith ContextNo ContextBoth Incorrect
@@ 4.2 Information Potential Across Corpora
[…]
Wikipedia Datasets (IP: 0.110–0.180). Both EPFL-related and Venice-related Wikipedia subsets show moderate IP with smaller WC–NC gaps (WC $\sim 97\%$ vs. NC $\sim 82$$84\%$). The attenuated gain indicates substantial overlap with pretraining data or readily retrievable priors, aligning with their broad web availability (Fig. 2).
[…]
@@ 4.3 Comparison between open and closed source model
Both Llama 3 70B and GPT-4o show similar dataset rankings for information potential: EPFL manuscripts consistently score about twice as high as Wikipedia. GPT-4o yields slightly lower IP than Llama 70B, suggesting stronger baseline knowledge retention, but both models reach near-perfect accuracy ($97\%$) with context, confirming robust comprehension when relevant text is provided.
A broader survey of models could clarify how architecture and scale shape IP, but here our focus is on validating it as a general metric across different model families.
5 Conclusion
This work introduced a lightweight pipeline to estimate the information potential (IP) of text collections for LLMs without retraining or fine-tuning. Our contributions are: